Good Time Company is a regional chain department store. It will remain in business for one more year. The probability of a boom year is 60 percent and the probability of a recession is 40 percent. It is projected that the company will generate a total cash flow of $203 million in a boom year and $94 million in a recession. The company's required debt payment at the end of the year is $128 million. The market value of the company’s outstanding debt is $101 million. The company pays no taxes.

  

a.

What payoff do bondholders expect to receive in the event of a recession? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, rounded to the nearest whole number, e.g., 1,234,567.)

b. What is the promised return on the company's debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
c. What is the expected return on the company's debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a). The expected payoff to bondholders is the face value of debt or the value of the company, whichever is less. Since the value of the company in a recession is $94 million and the required debt payment in one year is $128 million, bondholders will receive the lesser amount, or $94 million.

b). Promised Return = [Face Value of Debt / Market Value of Debt] - 1

= [$128,000,000 / $101,000,000] - 1 = 1.2673 - 1 = 0.2673, or 26.73%

c). Expected payment to bondholders = [0.60 * $128,000,000] + [0.40 * $94,000,000]

= $76,800,000 + $37,600,000 = $114,400,000

Expected return = (Expected value of debt / Market value of debt) - 1

= ($114,400,000 / $101,000,000] - 1 = 1.1327 - 1 = 0.1327, or 13.27%